OpenStudy (anonymous):

Describe a function g(x) in terms of f(x) if the graph of g is obtained by reflecting the graph of f about the x-axis and if it is horizontally stretched by a factor of 7 when compared to the graph of f. g(x) = Af(Bx) + C where A=? B=? C=?

OpenStudy (anonymous):

g(x) = -f((1/7)x) A = -1, B = 1/7, C = 0?

OpenStudy (anonymous):

correct. thanks ^^